# DynFOA: Generating First-Order Ambisonics with Conditional Diffusion for Dynamic and Acoustically Complex 360-Degree Videos

### v1

@@ 2.5.3 Multi-Condition Encoder and Cross-Modal Fusion
All conditional features are projected into a shared latent space before being injected into the diffusion U-Net. Modulation layers and cross-attention mechanisms fuse geometry, material, and propagation cues with descriptors of dynamic and multiple sound sources [wang2024multicond]. This cross-modal integration guides the denoising trajectory, ensuring that the generated FOA respects physical propagation constraints while maintaining perceptual consistency across time and sources.
@@ 2.5.4 Runtime Rendering and Head-Tracking
At inference, the diffusion model generates FOA conditioned on the reconstructed scene and dynamic context. The synthesized FOA signals are rotated according to the listener’s head orientation to maintain spatial alignment under head tracking [zhang2024audiovisual, li2023foa]. Finally, FOA are rendered to binaural signals using head-related transfer functions (HRTFs). This runtime process produces immersive spatial audio that adapts seamlessly to listener movement and complex multi-source environments.
@@ 3.1.1 Dataset Usage and Construction
Our experiment is conducted based on the YT360 and Sphere360 datasets [liu2025omniaudio], and uses a bilingual Chinese and English keyword semantic filtering algorithm combined with a manual review mechanism to construct our Dyn360 dataset. This dataset contains 600 strictly screened high-quality video clips that have been optimized to accurately reconstruct spatial audio in complex acoustic environments. All samples are normalized to 10 seconds in length, encoded in H.264 format, with a resolution of at least 720p and a stable frame rate of 30fps. The audio data is stored at a 16kHz sampling rate, 16-bit depth, and 4-channel format to generate high-quality FOA. All of the 600 clips are used for our quantitative evaluation reported later in Sec. 3.3.
@@ 3.1.2 Evaluation Metrics
Spatial accuracy evaluates the directional accuracy of the generated FOA using Direction of Arrival (DOA) estimation [zou2024doaeval], which measures the angular accuracy of the predicted sound source positions relative to the true source positions in 3D space.
Acoustic fidelity assesses different model’s ability to capture complex acoustic environments, using Signal-to-Noise Ratio (SNR) [chen2024snr] to measure audio clarity in the presence of background noise, and Early Decay Time (EDT) [li2024edt] to assess reverberation characteristics in various acoustic spaces, following established practices in spatial audio evaluation.
Audio distribution matching evaluates the similarity of feature distributions between real and generated FOA under our DynFOA and baselines. According to existing experimental progress, we compute the Fréchet Distance (FD) [cheng2025mmaudio, liu2025omniaudio] using a specialized spatial audio feature extractor. Furthermore, we introduced Short-Time Fourier Transform Error (STFT) [zhang2024stft] to measure spectral reconstruction accuracy, Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) [huang2024sisdr] to assess signal separation quality, and Kullback-Leibler (KL) [liu2025omniaudio] divergence to evaluate the statistical similarity between generated and reference audio distributions.
User study protocol was designed to conduct subjective evaluation using Mean Opinion Score (MOS-SQ) and Audio-Visual Alignment Fidelity (MOS-AF) [liu2025omniaudio] to assess the subjective user experience. Human evaluators rated the realism and synchronization of the generated FOA within the complex scenes including occlusion, reflections and reverberation.
@@ 3.1.3 Baselines
Since our work focuses on extracting scene information through 3D Gaussian reconstruction techniques and diffusion models to generate high-quality FOA from 360-degree videos, we construct the following benchmarks for comprehensive comparison: (1) Diff-SAGe [kushwaha2024diffsage], a diffusion-based spatial audio generation model that uses a diffusion process for FOA audio synthesis. (2) MMAudio [cheng2025mmaudio] with spatialization adaptation, this combined method integrates the state-of-the-art multimodal audio generation model MMAudio [cheng2025mmaudio] and an audio spatialization component. The spatialization component uses spatial angle estimation to convert the generated mono/stereo audio into a first-order surround sound FOA format. (3) OmniAudio [liu2025omniaudio], the original OmniAudio framework for converting 360-degree videos into spatial audio, which is our main benchmark for spatial audio synthesis (without using 3D Gaussian scene reconstruction). (4) ViSAGe [kim2025visage] , a traditional model designed to generate spatial audio from video input, focusing on the spatial audio generation task and performing well on directional audio synthesis. To ensure fair and consistent comparisons, we train and test all the models using proportional splits based on the Dyn360 dataset. Each baseline model takes a panoramic 360-degree video as its primary input, while the combined model accepts optional text prompts when available.

### v2

@@ 2.5.3 Multi-Condition Encoder and Cross-Modal Fusion
All conditional features are projected into a shared latent space before being injected into the diffusion U-Net. Modulation layers and cross-attention mechanisms fuse geometry, material, and propagation cues with descriptors of dynamic and multiple sound sources [rombach2022high]. This cross-modal integration guides the denoising trajectory, ensuring that the generated FOA respects physical propagation constraints while maintaining perceptual consistency across time and sources.
@@ 2.5.4 Runtime Rendering and Head-Tracking
At inference, the diffusion model generates FOA conditioned on the reconstructed scene and dynamic context. The synthesized FOA signals are rotated according to the listener’s head orientation to maintain spatial alignment under head tracking [morgado2020learning, zotter2019ambisonics]. Finally, FOA are rendered to binaural signals using head-related transfer functions (HRTFs). This runtime process produces immersive spatial audio that adapts seamlessly to listener movement and complex multi-source environments.
@@ 3.1.1 Dataset Usage and Construction
Our experiment is conducted based on the YT360 and Sphere360 datasets [liuomniaudio], and uses a bilingual Chinese and English keyword semantic filtering algorithm combined with a manual review mechanism to construct our Dyn360 dataset. This dataset contains 600 strictly screened high-quality video clips that have been optimized to accurately reconstruct spatial audio in complex acoustic environments. All samples are normalized to 10 seconds in length, encoded in H.264 format, with a resolution of at least 720p and a stable frame rate of 30fps. The audio data is stored at a 16kHz sampling rate, 16-bit depth, and 4-channel format to generate high-quality FOA. All of the 600 clips are used for our quantitative evaluation reported later in Sec. 3.3.
@@ 3.1.2 Evaluation Metrics
Spatial accuracy evaluates the directional accuracy of the generated FOA using Direction of Arrival (DOA) estimation [chen2010introduction], which measures the angular accuracy of the predicted sound source positions relative to the true source positions in 3D space.
Acoustic fidelity assesses different model’s ability to capture complex acoustic environments, using Signal-to-Noise Ratio (SNR) [loizou2007speech] to measure audio clarity in the presence of background noise, and Early Decay Time (EDT) [cerda2009room] to assess reverberation characteristics in various acoustic spaces, following established practices in spatial audio evaluation.
Audio distribution matching evaluates the similarity of feature distributions between real and generated FOA under our DynFOA and baselines. According to existing experimental progress, we compute the Fréchet Distance (FD) [cheng2025mmaudio, liuomniaudio] using a specialized spatial audio feature extractor. Furthermore, we introduced Short-Time Fourier Transform Error (STFT) [yamamoto2020parallel] to measure spectral reconstruction accuracy, Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) [jepsen2025study] to assess signal separation quality, and Kullback-Leibler (KL) [liuomniaudio] divergence to evaluate the statistical similarity between generated and reference audio distributions.
User study protocol was designed to conduct subjective evaluation using Mean Opinion Score (MOS-SQ) and Audio-Visual Alignment Fidelity (MOS-AF) [liuomniaudio] to assess the subjective user experience. Human evaluators rated the realism and synchronization of the generated FOA within the complex scenes including occlusion, reflections and reverberation.
@@ 3.1.3 Baselines
Since our work focuses on extracting scene information through 3D Gaussian reconstruction techniques and diffusion models to generate high-quality FOA from 360-degree videos, we construct the following benchmarks for comprehensive comparison: (1) Diff-SAGe [kushwaha2025diff], a diffusion-based spatial audio generation model that uses a diffusion process for FOA audio synthesis. (2) MMAudio [cheng2025mmaudio] with spatialization adaptation, this combined method integrates the state-of-the-art multimodal audio generation model MMAudio [cheng2025mmaudio] and an audio spatialization component. The spatialization component uses spatial angle estimation to convert the generated mono/stereo audio into a first-order surround sound FOA format. (3) OmniAudio [liuomniaudio], the original OmniAudio framework for converting 360-degree videos into spatial audio, which is our main benchmark for spatial audio synthesis (without using 3D Gaussian scene reconstruction). (4) ViSAGe [kim2025visage] , a traditional model designed to generate spatial audio from video input, focusing on the spatial audio generation task and performing well on directional audio synthesis. To ensure fair and consistent comparisons, we train and test all the models using proportional splits based on the Dyn360 dataset. Each baseline model takes a panoramic 360-degree video as its primary input, while the combined model accepts optional text prompts when available.